OpenStudy (anonymous):

2 2/3 divided by 1 8/9?? simplified form??

OpenStudy (anonymous):

2 and 2/3 = 8/3 1 and 8/9 = 17/9 8/3 divided by 17/9 = 8/3 * 9/17 = 72 / 51

OpenStudy (anonymous):

has to be a mixed number